what is this problem and put it in standard form:-5y=-3/4-4
Answer by checkley71(8403)

-5Y=-3/4-4
-5Y=-3 3/4
-5Y=-4.75
Y=4.75/5
Y=.95
PROOF
-5*.95=-.75-4
-4.75=-4.75